OpenStudy (anonymous):

Is the equation P = 500(1.03)n a model of growth or exponential decay? A. Growth B. Decay What is the value of f(x) = 20 (2x) when x = 3? A. f(3) = 120 B. f(3) = 64,000 C. f(3) = 160 D. f(3) = 80,000

OpenStudy (ghazi):

growth because exponent is positive

OpenStudy (ghazi):

when f(x)=20(2x) , just plug in the value put x=3 f(3) = 20(2*3)=20*6

OpenStudy (anonymous):

So f(3) = 120 ??

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

correct :)

OpenStudy (anonymous):

thank you :)

OpenStudy (ghazi):

:D